A student is standing at the top of a 10m tall building when they kick a soccer ball off the roof at a speed of 15m/s, at an angle of 30o above the horizontal. The ball then lands on the ground, some distance from the building.


1
Expert's answer
2021-09-24T09:25:24-0400

"g =9.8\frac{m}{s^2}"

"x_0=0; y_0=10"

"\vec v= 15\frac{m}{s}"

"\alpha=30\degree"

"\vec v= \vec v_x+\vec v_y"

"\vec v_x=\vec v\cos\alpha\approx13\frac{m}{s}"

"\vec v_y=\vec v\sin\alpha=7.5\frac{m}{s}"

"y =y_0+v_yt-\frac{gt^2}{2}"

"\text{When the ball falls to the ground:}"

"y=0"

"y_0+v_yt-\frac{gt^2}{2}=0"

"10+7.5t -4.9t^2=0"

"t =2.39s"

"x = x_0+\vec v_xt= 13*2.39=31.07 m"

"\text{Answer: distance from the building to where the ball lands is }31.07m"